OpenStudy (anonymous):

what the derivative for the function y=2.2^x ? if you can plzz

OpenStudy (anonymous):

y'=x2.2^x-1

OpenStudy (anonymous):

\[y = 2.2^x \Rightarrow \ln(y) = \ln(2.2^x) \Rightarrow \ln(y) = xln(2.2)\] Taking the derivative of both sides we have: \[\frac{1}{y}y' = \ln(2.2) \Rightarrow y' = yln(2.2)\] but: \[y = 2.2^x\] so plugging that in we obtain our answer: \[y' = \ln(2.2)*2.2^x\]

OpenStudy (anonymous):

hey why do you put ln you can easily solve it using the basic differentiation rules

OpenStudy (anonymous):

He is explaining that y = a^x means x= log_a y.

OpenStudy (anonymous):

@hashir your answer is wrong. you cant treat: \[y = a^x\] like: \[y = x^n\] They are totally different.